# PUMA: Efficient Continual Graph Learning with Graph Condensation

### v1

@@ 5.2 Overall Results
The condensation-based CGL methods are compared with all baselines in both class-IL and task-IL settings. AP is used to evaluate the average model performance of all learned tasks at the end of the task streaming, and BWT (also known as average forgetting (AF)) implies the forgetting problem of the model during continual learning. Table II shows the overall performance of all baselines and the PUMA in the class-IL CGL setting. CaT achieves the state-of-the-art performance compared with all other CGL baselines and can match the ideal Joint performance in the Arxiv, Reddit and Products by only maintaining a synthetic memory bank whose budget ratio is only 0.01. Besides, the results show that condensation-based memory bank has a smaller BWT, which means condensation cannot only preserve the historical knowledge of the model but reduce the negative effects on the previous tasks while training the current task to alleviate the catastrophic forgetting problem. PUMA and CaT outperforms other baselines in CoraFull but does not reach the Joint performance with two potential reasons: (1) the 0.01 budget ratio for CoraFull limits the replayed graph to only four nodes, which is extremely small to contain sufficient information; (2) CoraFull has 35 tasks, which is more than other datasets and difficult to retain historical knowledge.
Other baselines can hardly match the performance of condensation-based memory bank. Finetuning is easy to forget the previous knowledge since it only uses the newly incoming graph to update the model. Regularisation-based methods (e.g., EWC, MAS, GEM, TWP) also have unsatisfactory performance since adding overhead restrictions to the model will lead to bad model plasticity during the long streaming tasks. As a distillation method, LwF hardly handles the class-IL setting in the CGL. ER-GNN does not have reasonable results in all benchmarks for the sampling-based replay methods since there is a severe imbalanced training problem. SSM stores sparsified subgraphs in the memory bank, which can preserve the topological information for the historical graph data. Although SSM has a good performance, it still has a gap to Joint or condensation-based methods.
It is worth noting that even though all historical data can be used for training, Joint also has a negative BWT. The reason is that the class-IL setting requires the model to increase the output layer dimension as the new classes emerge, where the model cannot perfectly remember all previous knowledge. The results of HPNs for the class-IL are not provided since it is not designed for the class-IL setting.
Table III shows the overall performance under the task-IL setting. Compared to the class-IL setting, task-IL is much easier, and all baseline methods get reasonable results. The CaT can match the Joint method with a 0.01 budget ratio.
@@ 5.3 Ablation Study
The PUMA framework has two key components, pseudo-label guided memory bank and retraining. To study their effectiveness, different variants are evaluated, and mAP of these variants are reported in Table IV to indicate the overall performance. The variant without PL denotes only labelled nodes are used in the condensation process, and the variant without Re indicates the CGL model updates their weights from the learned knowledge of previous tasks. According to Table IV, compared to the variant without both components, the variant using PL improves both AP and mAP in the CoraFull, Arxiv and Products datasets but keeps the same in the Reddit dataset. The main reason is that the information of labelled data in CoraFull, Arxiv and Products are not sufficiently restore the node feature distribution, but the information of labelled nodes in the Reddit dataset is informative enough. Compared to the variant without retraining, the variant with retraining improves the overall performance, especially on small datasets (e.g., CoraFull and Arxiv).
@@
5.4 Effectiveness and Efficiency of Condensation-based Memory Banks
To analyse the effectiveness and efficiency of condensation-based memory banks (e.g., CaT and PUMA), four typical memory banks are evaluated with four budget ratios, i.e., 0.005, 0.01, 0.05, and 0.1. Specifically, 0.005 is an extremely limited budget ratio, under which the memory bank only contains 2-node replayed graphs on the CoraFull dataset. While 0.1 is a large budget ratio, which represents the size of the memory bank equals to 10% of the size of the entire training set. For a fair comparison with PUMA, the TiM and retraining is applied for ER-GNN, SSM and CaT.
@@ 5.4 Effectiveness and Efficiency of Condensation-based Memory Banks
5.4.1 Different Memory Banks
Fig. 3 demonstrates that the PUMA is more effective than existing sampling-based memory banks. PUMA converges to optimal performance much quicker. PUMA gets the best performance in all evaluated cases. PUMA significantly outperforms other sampling-based memory banks when the budget ratio is relatively small (e.g., 0.005, 0.01). Besides, condensation-based memory banks have small standard errors at the limited memory budgets, demonstrating the robustness.
@@ 5.4.2 Budget Efficiency
The advantage of the condensed graph is to keep the information of the original graph while reducing the graph size significantly. Fig. 3 shows that the CaT and PUMA outperform the sampling-based methods in achieving higher performance within a more limited budget. In all datasets, the sampling-based methods have a huge performance gap to condensation-based methods. For example, PUMA with 0.005-budget ratio can outperform or match the 0.1-budget ratio sampling-based methods on CoraFull, Reddit and Products.
On the one hand, CaT and PUMA use less memory spaces to accurately approximate the historical data distribution. On the other hand, in the training phase, the model needs to propagate messages in the memory bank. Therefore, a small memory bank can improve both storage and computation efficiency.
@@ 5.5 Balanced Learning with TiM
5.5.1 Different Methods with TiM
TiM is a plug-and-play training scheme for all existing replay-based CGL methods. Table  V shows the mAP for different replay-based CGL methods with and without TiM. It clearly shows that the TiM scheme can improve the overall average performance with all memory bank generation methods. The reason is that the TiM can ensure training graphs for the CGL models have a similar size to deal with the imbalanced issue, which can solve the catastrophic forgetting problem.
TABLE V: Comparison of mAP (%) for replay-based CGL methods without and with TiM.
@@ 5.5.2 Visualisation
The performance matrices of ER-GNN, SSM, CGM, and those memory banks training with TiM (i.e., EaT, SaT can CaT) on the CoraFull, Arxiv, Reddit and Products datasets under 0.01 budget ratio are visualised in Fig. 4. All memory banks without TiM struggle with remembering the previous knowledge since the scale gap between the newly incoming graph and replayed graphs in the memory bank. After using the TiM scheme, the performance matrices show the forgetting process slows down (i.e., the colour of each column is not changed a lot), which indicates the catastrophic forgetting problem is alleviated as the imbalanced training issue is tackled.
@@ 5.6 Effectiveness of Retraining
Although TiM effectively mitigates the issue of class training sample imbalance, it introduces a new challenge during continual training: imbalance in task losses. This occurs because memories added earlier to the memory bank have been sufficiently learned, resulting in smaller losses compared to those added later. For alleviating this problem, retraining strategy is proposed to train the CGL model in the memories from scratch. This section will discuss the effectiveness of retraining strategy by applying it to different memory banks.
TABLE VI: Comparison of mAP (%) for replay-based CGL methods without and with retraining.
@@ 5.6.1 Effectiveness
Table VI represents the mAP of different replay-based methods without and with retraining. The metric mAP is selected for comparative analysis, as it enables a comprehensive evaluation of accuracy at each step of the continual learning process, rather than focusing solely on the final step performance.
In the presented data, the efficacy of retraining in enhancing the performance of replay-based CGL methods is clearly demonstrated across various datasets. In the context of continual training, this aspect predisposes the model to overfitting on previous memories. Furthermore, the issue of imbalanced loss exacerbates this tendency, as it hinders the model’s ability to effectively learn the decision boundaries between newly added memories and historical ones. Consequently, retraining becomes crucial in this scenario, as it helps in recalibrating the model’s knowledge and adaptability to both new and existing data, thereby improving its overall performance.
@@ 5.6.2 Visualisation
Fig. 5 shows the accuracy changes of the fist task after receiving the last task of the graph streaming during continual training. For the CoraFull, Arxiv and Products datasets, the accuracy without retraining has a shake and converge then. Although the retrained model needs more steps to converge, the accuracy with retraining can outperform the continual training. For the Reddit dataset, retrain the CGL model cannot benefit a lot since the model accuracy is already high.
Figure 5: The accuracy (%) changes of the first task during continual training of the last task on CoraFull, Arxiv, Reddit, and Products datasets.
@@ 5.7 Condense More by Pseudo-Labelling
The impact of pseudo-labelling on the accuracy of replay-based CGL methods, such as ER-GNN, SSM, CaT and PUMA, is evaluated in this study. The comparison involves scenarios with and without the integration of pseudo labels during the memorisation stage. Table VII presents the mAP for different CGL methods under both conditions.
It is observed that for condensation-based graph memory, incorporating pseudo-labeling proves to be effective. This approach aids in training a robust classifier, which in turn generates useful pseudo-labels. Additionally, introducing pseudo labels into incoming graphs expands the training dataset, providing a more comprehensive set of supervisory signals. Such an expansion significantly enhances the generalisation capabilities of the CGL model, leading to improved accuracy in inferring labels of test nodes.
However, it is noted that the pseudo-labelling technique does not yield similar benefits for current sampling-based memory banks. The limitation primarily arises due to the reliance of pseudo-labelling classifiers on memory banks for training. Sampling-based memory banks with limited budgets might not contain sufficient information, which hampers the development of an effective classifier. On the other hand, the fixed budgets of sampling-based memory banks restrict their ability to add more information. This is in contrast to condensation-based methods, which can incorporate additional information despite similar budget constraints by discovering more unlabelled nodes.
TABLE VII: mAP of replay-based CGL methods with and without pseudo-labelling
@@ 5.8 Wide Graph Encoder
A wide graph encoder includes more neurons with random initialised weights which randomly extracts features with non-linearity. Fig. 6 shows reducing the distance between original and condensed graph embeddings generated by a narrow encoders is not sufficient since there is still a clear distribution gap between them once the encoder is reinitialised. This gap is decreased by wider graph encoders.
The more neurons that can be obtained at one time, the clearer the potential transformations of the original data in the initialisation space will be, and the easier it will be for fitting the distribution of data through different networks. It is necessary to have an accurate distribution of data because in continual learning, new classes emerge continuously, and during the replay process, the model needs to relearn the decision boundary between different classes. However, more neurons spends more computation resources, multiple random encoders are used in practice.
@@
Figure 6: Kernel density estimation (KDE) plot of mean embeddings. It compares KDE of mean embeddings at different stages: ‘Target’ (original graph’s embedding), ‘Target-Reinit’ (original graph’s embedding by a reinitialised encoder), and ‘PUMA-Reinit’ (condensed graph’s embedding by a reinitialised encoder). The encoder dimensions of left figure and right figure are 800 and 12800 respectively.
@@ 5.9 Parameter Sensitivity
There are several hyperparameters in condensed memory bank generation, including the budget ratio for the replayed graph, which is already evaluated in Fig. 3. This section will discuss the choices of dimension of graph encoders and usage of activation during encoding.
@@ 5.9.1 Different Dimensional Graph Encoders
This experiment explores the sufficient dimensions of graph encoders to optimise the MMD loss. The condense iterations remain 500, and the graph encoder only randomly initialise once before condensation. Table 7 shows the AP (%) of the model trained by the PUMA generated by various dimensional graph encoders after handling each incoming task during the continual learning process.
Figure 7: The average performance (AP) changes in the whole continual learning process with memory banks generated by different dimensional graph encoders. Except for CoraFull dataset, all other datasets only update the condensed graph 500 times under only one random parameterised graph encoder. For CoraFull, 50 random graph encoders are used to update the condensed graph 100 times individually for reducing the distribution distance in a larger embedding space since CoraFull’s 8710-dimensional features imply a larger model parameter initialisation space.
@@ 5.9.2 Neuron Activation
This experiment will compare the use of activation functions in encoders for graph encoding. The standard node classification for CGL will still utilise an encoder with an activation function. mAP is employed to measure effectiveness. Table VIII illustrates that for varying budget ratios, encoders with activation functions are more competitive options for encoding graphs overall.
TABLE VIII: mAP (%) of different graph encoders with varying budget ratios
@@ 5.10 Time Efficiency
For replay-based CGL methods, memory generation and model training are two main parts for requiring computation resources. Fig. 8 compares the total time (memorisation plus training time) for ER-GNN, SSM, CaT and PUMA methods. Experiments are ran on CoraFull and Products datasets.
Figure 8: Total time (i.e., memorisation time plus training time) for different replay-based CGL methods in the streaming of Arxiv and Products datasets. x-axis represents total time (s) and y-axis indicates the AP (%). The size of scatter implies the budget of the memory bank.
For sampling-based methods (e.g., ER-GNN and SSM), memory generation process is efficient. Although the time of graph condensation cannot be ignored as sampling-based methods, the model accuracy are much better than sampling-based methods. For CaT, the main computation cost for the memory generation is the feature aggregation in the original graph. Since the condensed graph is significantly smaller than the original graph, the time of feature aggregation in the condensed graph can be ignored. Because the original graph is fixed, once the original graph’s aggregated features are obtained, it can be stored in the memory for reduce the repeatedly computation. Once weights of the graph encoder is reinitialised, the embedding of original graph in this embedding space is needed to recalculate. This step can be reduced by increasing the dimension of graph encoders, and it can even be reduced to 1 when the dimension of graph encoder is sufficient.
The memory bank of PUMA is edge-free, the feature aggregation operation in the training phase can be ignored, weights of each layer can be learned by a MLP model. The performance of PUMA can match the Joint method, but the operation time is faster.
Although edge-free condensed graph can accelerate the training process, it can not counteract the increasing time in the memory generation process. Considering the accuracy improvement of PUMA, it is still competitive.

### v2

@@ 5.2 Overall Results
The CaT and PUMA are compared with all baselines in both class-IL and task-IL settings. AP is used to evaluate the average model performance of all learned tasks at the end of the task streaming, and BWT denotes the forgetting problem during continual learning. Table II shows the overall performance of all methods in the class-IL CGL setting. CaT and PUMA achieve state-of-the-art performance compared with all other CGL baselines and can match the ideal Joint performance in the Arxiv, Reddit and Products by only maintaining a synthetic memory bank with a 0.005 budget ratio. Besides, the results show that a condensation-based memory bank has a smaller BWT, which means that the condensation not only preserves the historical knowledge but reduces the negative effects on the previous tasks while training the current task to alleviate the catastrophic forgetting problem. PUMA and CaT outperform other baselines in CoraFull but do not reach the Joint performance with two potential reasons: (1) the 0.005 budget ratio for CoraFull limits the replayed graph to only two nodes, which is extremely small to contain sufficient information; (2) CoraFull has 35 tasks, which is more than other datasets and difficult to retain historical knowledge. The same issue of limited node sizes occurs in both the A-Computer dataset (8 nodes per replayed graph) and the A-Photo dataset (6 nodes per replayed graph), which leads to a little performance gap when compared with the joint method.
Other baselines can hardly match the performance of condensation-based memory bank. Finetuning is easy to forget the previous knowledge since it only uses the newly incoming graph to update the model. Regularisation-based methods (e.g., EWC, MAS, GEM and TWP) also have unsatisfactory performance since adding overhead restrictions to the model will lead to decreased model plasticity during the long streaming of tasks. As a distillation method, LwF hardly handles the class-IL setting in the CGL. ER-GNN does not have reasonable results in all benchmarks for the sampling-based replay methods since there is a severe imbalanced training problem. SSM stores sparsified subgraphs in the memory bank, which can preserve the topological information for the historical graph data. Although SSM has a good performance, it still has a gap to Joint or condensation-based methods. DSLR enhances the diversity of the memory bank but does not significantly improve model performance due to budget limitations and imbalance training issues. Ensemble methods that combine the memory bank with the regularisation techniques (e.g., MetaCLGraph and ER-LS-GS) still suffer inadequate historical information and imbalanced training problems.
Table III shows the overall performance under the task-IL setting. Compared to class-IL, task-IL is much easier, and all baselines achieve a reasonable result. The CaT and PUMA can match the Joint method with a 0.005 budget ratio.
@@ 5.3 Ablation Study
The PUMA has two key components, pseudo-label guided memory bank and retraining. To study the effectiveness, different variants are evaluated, and mAP of these variants is reported in Table IV to indicate overall performance. The variant without PL indicates that only labelled nodes are used in the condensation process, while the variant without Re implies that the CGL model updates its weights based on the learned knowledge from previous tasks. According to Table IV, compared to the variant lacking both components, the variant using PL improves the performance in the CoraFull, Arxiv, and Products datasets, but shows no change in the Reddit dataset. This is because the labelled data in CoraFull, Arxiv, and Products do not sufficiently capture the node feature distribution, whereas the labelled nodes in the Reddit dataset are sufficiently informative. Additionally, the variant with retraining enhances overall performance, particularly in smaller datasets such as CoraFull and Arxiv.
@@
5.4 Efficacy of Condensation-based Memory Banks
To analyse the Efficacy of condensation-based memory banks (e.g., CaT and PUMA), four memory banks are evaluated with four budget ratios, i.e., 0.005, 0.01, 0.05, and 0.1. For a fair comparison with PUMA, the TiM and retraining are applied for ER-GNN, SSM and CaT.
@@ 5.4.1 Budget Efficiency
The advantage of the condensed graph is to keep the information of the original graph while reducing the graph size. Fig. 3 shows that the CaT and PUMA outperform the sampling-based methods with a more limited budget. In all scenarios, the sampling-based methods have a huge performance gap to CaT and PUMA. PUMA with 0.005-budget ratio can outperform or match the 0.1-budget ratio sampling-based methods on CoraFull, Reddit and Products.
On one hand, CaT and PUMA use less memory space to approximate the historical data. On the other, in the training phase, a smaller memory bank can benefit the model with storage and computation efficiency during message passing.
5.4.2 Label Distribution of Memory Bank
Although TiM ensures that the training graphs have a balanced size, the number of nodes in each class remains proportional to the incoming graphs in the default setting. Consequently, if the incoming graph exhibits imbalanced class sizes, this imbalance will be reflected in the training graphs. To investigate how label distribution impacts condensed graph memory, we studied the effects of condensing the incoming graph with a balanced label distribution. Table V presents the performance of CaT and PUMA on the Arxiv dataset using different label distribution strategies, specifically proportional and balanced distributions. The results demonstrate that a balanced label distribution results in a performance decline. This decline occurs because the condensation ratio increases for the larger classes while it relaxes for the smaller classes compared to the proportional label distribution. This finding suggests that larger classes, which contain more information, benefit from being allocated a larger number of nodes.
TABLE V: CaT and PUMA with different label initialisation strategies on Arxiv. Proportional label distribution means each replayed graph has 30 nodes, with the size of each class proportional to the incoming graph. In contrast, balanced label distribution ensures each class has 15 nodes.
Method Label Distribution AP BWT
CaT Proportional 65.4±1.9 -12.4±1.5
Balanced 63.6±1.3 -13.0±1.1
PUMA Proportional 68.4±0.2 -10.9±0.4
Balanced 66.5±0.5 -10.8±0.7
@@ 5.5 Balanced Learning with TiM
5.5.1 Replay-based CGL Methods with TiM
TiM is a plug-and-play training scheme for replay-based CGL methods. Table  VI shows the mAP for different replay-based CGL methods with and without TiM. It demonstrates that the TiM scheme can improve the average performance with all memory bank generation methods. The reason is that the TiM can ensure training graphs for the CGL models have a similar size to deal with the imbalanced issue, which can mitigate the catastrophic forgetting problem.
TABLE VI: Comparison of mAP (%) for replay-based CGL methods without and with TiM.
@@ 5.5.2 Visualisation
The performance matrices of PUMA on the CoraFull, Arxiv, Reddit and Products datasets under 0.01 budget ratio are visualised in Fig. 4. All memory banks without TiM struggle with remembering the previous knowledge since the scale gap between the newly incoming graph and replayed graphs in the memory bank. After using the TiM scheme, the performance matrices show the forgetting process slows down (i.e., the colour of each column is not changed a lot), which indicates the catastrophic forgetting problem is alleviated as the imbalanced training issue is tackled.
@@ 5.6 Effectiveness of Retraining
Although TiM effectively mitigates the issue of training graph imbalance, it introduces a new challenge during continual training: imbalance in task losses. Memories added earlier to the memory bank are sufficiently learned, resulting in smaller losses compared to those added later. To alleviate this problem, a retraining strategy is proposed to train the model in the memories from scratch.
TABLE VII: Comparison of mAP (%) for replay-based CGL methods without and with retraining.
@@ 5.6.1 Effectiveness
Table VII represents the mAP of different replay-based methods without and with retraining. The metric mAP is selected as it enables a comprehensive evaluation of accuracy at each step of the continual learning process, rather than focusing solely on the final step performance.
In the experiment, the efficacy of retraining for replay-based CGL methods is clearly demonstrated across various datasets. During continual training, models are prone to overfitting on previous memories. Additionally, the issue of imbalanced loss exacerbates this tendency, hindering the model’s ability to effectively learn the decision boundaries between newly added and historical memories. Consequently, retraining becomes crucial, as it recalibrates the model’s knowledge and adaptability to both new and existing data, thereby improving overall performance.
@@ 5.6.2 Visualisation
Fig. 5 shows the accuracy changes of the first task after receiving the last task of the graph streaming during continual training. For the CoraFull, Arxiv and Products datasets, the accuracy without retraining has a shake and converge then. Although the retrained model needs more steps to converge, the accuracy with retraining can outperform the continual training. For the Reddit dataset, retraining the CGL model cannot benefit a lot since the model accuracy is already high.
Figure 5: The accuracy changes of the first task when training with the last task on four datasets.
@@ 5.7 Condense More by Pseudo-Labelling
The impact of pseudo-labelling on the accuracy of replay-based CGL methods, such as ER-GNN, SSM, CaT and PUMA, is evaluated in this study. The comparison involves scenarios with and without the integration of pseudo labels during the memorisation stage. Table VIII presents the mAP for different CGL methods under both conditions.
It is observed that for condensation-based graph memory, incorporating pseudo-labelling proves to be effective. This approach aids in training a robust classifier, which in turn generates useful pseudo-labels. Additionally, introducing pseudo labels into incoming graphs expands the training dataset, providing a more comprehensive set of supervisory signals. Such an expansion significantly enhances the generalisation capabilities of the CGL model, leading to improved accuracy in inferring labels of test nodes. However, it is noted that the pseudo-labelling technique does not yield similar benefits for current sampling-based memory banks. The limitation primarily arises due to the reliance of pseudo-labelling classifiers on memory banks for training.
TABLE VIII: mAP of replay-based CGL methods with and without pseudo-labelling.
@@ 5.8 Wide Graph Encoder
A wide graph encoder with randomly initialised weights can randomly extract features with non-linearity. As shown in Fig. 6, reducing the distance between the original and condensed graph embeddings generated by a narrow encoder is not sufficient, as there remains a clear distribution gap between them when the encoder is reinitialised and this gap is reduced by using wider graph encoders.
The more neurons are involved at one time, the clearer the potential transformations of the original data in the initialisation space will be, making it easier to fit the data distribution through different networks. Having an accurate data distribution is crucial in continual learning because new classes continuously emerge, and during the replay process, the model needs to relearn the decision boundary between different classes. However, using more neurons requires more computational resources, so multiple random encoders are used in practice.
@@
Figure 6: Kernel density estimation (KDE) plot of mean embeddings. It compares KDE of mean embeddings at different stages: ‘Target’ (original graph’s embedding), ‘Target-Reinit’ (original graph’s embedding by a reinitialised encoder), and ‘PUMA-Reinit’ (condensed graph’s embedding by a reinitialised encoder). The encoder dimensions of the left figure and right figure are 800 and 12800, respectively.
@@ 5.9 Parameter Sensitivity
There are several hyperparameters in condensed memory bank generation, including the budget ratio for the replayed graph, which is already evaluated in Fig. 3. This section will discuss the choices of the dimension of graph encoders and the usage of activation during encoding.
@@ 5.9.1 Different Dimensional Graph Encoders
This experiment explores the sufficient dimensions of graph encoders to optimise the MMD loss. The condense iterations remain 500, and the graph encoder is only randomly initialised once before condensation. Figure 7 shows the AP of the model trained by the PUMA generated by various dimensional graph encoders after handling each incoming task during the continual learning process.
Figure 7: The average performance (AP) changes in the whole continual learning process with memory banks generated by different dimensional graph encoders. Except for the CoraFull dataset, all other datasets only update the condensed graph 500 times under only one random parameterised graph encoder. For CoraFull, 50 random graph encoders are used to update the condensed graph 100 times individually for reducing the distribution distance in a larger embedding space since CoraFull’s 8710-dimensional features imply a larger model parameter initialisation space.
@@ 5.9.2 Neuron Activation
This experiment investigates the impact of using activation functions in encoders for CGL. The chosen activation function is ReLU. mAP is employed to measure effectiveness. Table IX illustrates that for varying datasets, encoders with activation functions are more competitive options for encoding graphs overall. It can be concluded that using ReLU can benefit the performance of CGL.
TABLE IX: mAP (%) of different graph encoders with varying budget ratios
@@ 5.10 Time Efficiency
For replay-based CGL methods, memory generation and model training are two main parts for requiring computation resources. Fig. 8 compares the total time (memorisation plus training time) for ER-GNN, SSM, CaT and PUMA methods on CoraFull and Products datasets.
Figure 8: Total time (i.e., memorisation time plus training time) for different replay-based CGL methods in the streaming of Arxiv and Products datasets. x-axis represents total time (s) and y-axis indicates the AP (%). The size of the scatter implies the budget of the memory bank.
For sampling-based methods (e.g., ER-GNN and SSM), memory generation process is efficient. Although the time of graph condensation cannot be ignored as sampling-based methods, the model accuracy is much better than sampling-based methods. For CaT and PUMA, the main computation cost for the memory generation is the feature aggregation in the original graph. Since the condensed graph is significantly smaller than the original graph, the time of feature aggregation in the condensed graph can be ignored.
With the help of an edge-free memory bank, PUMA can ignore the feature aggregation operation during the training phase, allowing the weights of each layer to be learned by an MLP model. As a result, PUMA’s performance can match that of the Joint method, but with faster operation times.